The set of all points that are one unit from (0,-1). Write the equation for the given set of points. Then graph your equation.
I can't figure out the equation and graph. Please help.
Thank you
Leona Anderson

Answer by scott8148(6628)

all points (in a plane) that are equidistant from a given point, form a circle

in this case, the center of the circle is (0, -1); and the radius is 1

for a circle centered at (h, k) with radius r, the general equation is (x-h)^2 + (y-k)^2 = r^2

in this case ___ (x-0)^2 + (y - -1)^2 = 1^2 ___ x^2 + y^2 + 2y + 1 = 1